# Simulating and investigating various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model

### v2

@@ II.2 States and Hamiltonian
$|d\rangle_{dist}|p\rangle_{prot}|n\rangle_{phn}$ (9)
where the first qubit $|d\rangle_{dist}$ is the relative position of two water molecules, which is detailed defined in Sec. II.1; the second quantum bit $|p\rangle_{prot}$ is the state corresponding to the proton, $p=0$ — proton is in the ground state $|\Phi_{0}^{pr}\rangle$ and $p=1$ — proton is in the excited state $|\Phi_{1}^{pr}\rangle$; the third qubit $|n\rangle_{phn}$ is the number of thermal phonon corresponding to the hydrogen bond. Here a special case of our model, where at most one phonon is pumped into the system (at this time, $n_{av}=1$), is proposed. When a hydrogen bond is formed, a phonon is emitted. When a hydrogen bond is broken, the phonon is absorbed.
@@
Table 1: Quantum states involving in evolution. The evolution involves a total of 7 states. For convenience, the following modifiers is used to define different states: “broken” means hydrogen bond is broken and two molecules are free; “critical” means the critical point of hydrogen bond formation and breaking is obtained; “stretched” means the proton of donor moves toward receptor due to the ion attraction and covalent bond is stretched; “excited” means proton state is excited; “ground” means proton state is ground; “stable” means hydrogen bond is stable. Especially, when the state is stable, it must also be stretched and ground. Besides states $|-1\rangle|0\rangle|1\rangle$ and $|-1\rangle|0\rangle|0\rangle$, the rest states can all defined as “unstable”. In addition, there are two cases for the ground state: with/without a phonon.
Figure 2: (online color) Quantum states involved in evolution. (a) $\sim$ (g) correspond to seven states for $\rm{H}_{2}\rm{O}$-related hydrogen bond model: $|2\rangle|1\rangle|0\rangle,\ |1\rangle|1\rangle|0\rangle,\ |0\rangle|1\rangle|0\rangle,\ |0\rangle|0\rangle|1\rangle,\ |0\rangle|0\rangle|0\rangle,\ |-1\rangle|0\rangle|1\rangle,\ |-1\rangle|0\rangle|0\rangle$. Here brown, red, green and purple circles represent oxygen atom, proton with excited state, proton with ground state and thermal phonon, respectively; solid and dashed black lines stand for ordinary and stretched covalent bond, respectively; ion attraction is denoted by gray dotted lines.
$H_{hb}$ is represented by the Hamiltonian of a system of two water molecules connected by hydrogen bond and has the following form
$H_{hb}=H_{dist}Cond(p=1)+H_{prot}Cond(d=0)$ (10)
where $H_{dist}$ describes proton tunneling energy and $H_{prot}$ describes Hamiltonian for transitions. $Cond(p=1)$ and $Cond(d=0)$ are conditional operators. The rotating wave approximation (RWA) is particularly useful in systems where the interactions are resonant or nearly resonant [53]. RWA is taken into account and the Hamiltonians is described in following form
@@
    $H_{dist}=\hbar\omega_{dist}\sigma^{{\dagger}}_{dist}\sigma_{dist}+g_{dist}\left(\sigma^{{\dagger}}_{dist}+\sigma_{dist}\right)$ (11)
    where $\omega_{dist}$ is frequency for tunneling of phonons; $g_{dist}$, which is located on the non-diagonal line of the Hamiltonian, is the strength for tunneling of proton. And definitions of operators $\sigma_{dist}$ and $\sigma^{{\dagger}}_{dist}$ are as follows
    $\sigma_{dist}|1\rangle_{dist}$ $=|0\rangle_{dist}$ (12)
    $\sigma_{dist}^{{\dagger}}|0\rangle_{dist}$ $=|1\rangle_{dist}$
@@
    $H_{prot}$ $=\hbar\omega_{phn}a^{{\dagger}}_{phn}a_{phn}+\hbar\omega_{prot}\sigma^{{\dagger}}_{prot}\sigma_{prot}$ (13)
    $+g_{prot}\left(a_{phn}\sigma^{{\dagger}}_{prot}+a^{{\dagger}}_{phn}\sigma_{prot}\right)$
    where $\omega_{phn}$ is phononic mode, $\omega_{prot}$ is mode for transitions of proton, and $\omega_{phn}=\omega_{prot}$; $g_{prot}$ is the strength for transitions of proton. And definitions of operators $a^{{\dagger}}_{phn}\sigma_{prot}$ and $a_{phn}\sigma_{prot}^{{\dagger}}$ are as follows
    $a^{{\dagger}}_{phn}\sigma_{prot}|1\rangle_{prot}|0\rangle_{phn}$ $=|0\rangle_{prot}|1\rangle_{phn}$ (14)
    $a_{phn}\sigma_{prot}^{{\dagger}}|0\rangle_{prot}|1\rangle_{phn}$ $=|1\rangle_{prot}|0\rangle_{phn}$
The complete quantum system ultimately constructed by the total Hamiltonian $H_{hb}$ can be intuitively represented in the form of a network, which is shown in Fig. 3. Interactions, dissipations and inflows between states can be intuitively seen in this figure, and the energy wandering of the whole system is determined clearly.
Figure 3: (online color) Network of quantum states and energy wandering. The quantum system consists of several states, between which there are some potential interactions, dissipations and inflows. In this figure, the dots represent the states and the edges — interactions, dissipations and inflows. Red dots stand for the states corresponding to the formation of stable hydrogen bond; green dot represents the state corresponding to the case that hydrogen bond is broken and two molecules are free; blue dots represent the states corresponding to unstable hydrogen bond. Black edges stand for the interactions ($g_{dist}$ and $g_{prot}$); purple edges represent the dissipations and inflows. In particular, $\gamma_{isol}$ includes both dissipation $\gamma_{isol}^{out}$ from state $|1\rangle|1\rangle|0\rangle$ to state $|2\rangle|1\rangle|0\rangle$ and inflow $\gamma_{isol}^{in}$ from state $|2\rangle|1\rangle|0\rangle$ to state $|1\rangle|1\rangle|0\rangle$.
@@ II.3 Quantum master equation
$A_{phn}|1\rangle_{phn}=|0\rangle_{phn}$ (15)
$A_{phn}^{{\dagger}}|0\rangle_{phn}=|1\rangle_{phn}$
And $\gamma_{phn}$ is the dissipation intensity for the escape of phonons from cavity to external environment. Similarly, the two quasi-decoherence factors, which are related to the qubit $|d\rangle_{dist}$ of the basic state, are defined as follows
$A_{bond}|0\rangle_{dist}=|-1\rangle_{dist}$ (16)
$A_{bond}^{{\dagger}}|-1\rangle_{dist}=|0\rangle_{dist}$
$A_{isol}|1\rangle_{dist}$ $=|2\rangle_{dist}$ (17)
$A_{isol}^{{\dagger}}|2\rangle_{dist}$ $=|1\rangle_{dist}$
@@
The QME, which is also called Lindbladian, is used to describe the time-dependent evolution of the density matrix of an open quantum system, which allow us to obtain the dependence of the probability of hydrogen bond formation on temperature. It can be seen as a natural result of the coupling between the system and the environment under the lowest-order perturbation and Markov approximation
$i\hbar\dot{\rho}=[H,\rho]+i\mathcal{L}(\rho)$ (18)
where $\mathcal{L}=L^{in}+L^{out}$. $L^{in}$ describes the inflow process and $L^{out}$ describes the dissipation process. Then
$L^{in}(\rho)=\gamma^{in}\left(A^{\dagger}\rho A-\frac{1}{2}\left(AA^{\dagger}\rho+\rho AA^{\dagger}\right)\right)$ (19)
$L^{out}(\rho)=\gamma^{out}\left(A\rho A^{\dagger}-\frac{1}{2}\left(A^{\dagger}A\rho+\rho A^{\dagger}A\right)\right)$ (20)
where $A$ and $A^{\dagger}$ can be replaced by $A_{bond},\ A_{isol},\ A_{phn},\ A_{bond}^{\dagger},\ A_{isol}^{\dagger}$ and $A_{phn}^{\dagger}$. The full form of the operator $\mathcal{L}$ with consideration of all types of inflows and dissipations is shown in Appx. A.
@@ III Numerical method
To solve the QME, various numerical methods such as the Euler method. The iteration consists of three steps. The first step is to calculate the unitary evolution of the density matrix
$\widetilde{\rho}(t+\Delta t)=e^{\frac{-iH\Delta t}{\hbar}}\rho(t)e^{\frac{iH\Delta t}{\hbar}}.$ (21)
where $\Delta t$ is the iteration time step and the second step is to apply the Lindblad superoperator (19) and (20) to the density matrix $\rho(t+\Delta t)$ at this moment
@@
The third step is to scale the density matrix to maintain its positive definiteness, Hermitian, normalization and other properties.
In addition to the Euler method, there is also the more accurate Runge-Kutta method. For quantum systems with particularly large states, the Monte Carlo wave function method [54] or the pure state vector method [55] can achieve higher computational efficiency. The $\rm{H}_{2}\rm{O}$-related hydrogen bond model contains a total of 7 states, so using the Euler method makes it easier for us to study the effects of various interactions and dissipations on evolution.
When solving the QME using the Euler method, The order of magnitude of the time step $\Delta t$ usually depends on the unit of the system’s Hamiltonian and the units of other related parameters. In optical Cavity QED experiments with high finesse optical resonators [56] or microwave resonators [57], the time step is determined by the measurement time, which is around 1 $\mu s$. However, for a single hydrogen bond, the time step reference measurement time of 1 $\mu s$ is not appropriate. Therefore, the characteristic time scale $\tau$ of the simulated dissipative quantum system is adopted, which can be estimated from the system energy and the reduced Planck constant
$\tau=\frac{\hbar}{E}$ (23)
where $E$ is the energy of the quantum system and the intensity factor $\gamma$ is usually smaller than $E$. When implementing the Euler method numerically, the time step $\Delta t$ should be significantly smaller than the characteristic time scale.
The energy of a single hydrogen bond $E_{H-bond}$ can be estimated from the average energy of hydrogen bonds in water and Avogadro’s constant, which is approximately $0.217655\ eV$. Using joule as the energy unit may cause machine error due to too small parameters, so $eV$ is used as the energy unit. And the characteristic time scale $\tau$ is about $3.175316\times 10^{-15}\ s$. The detailed calculation process is shown in Appx. B.
@@ IV Simulations and results
For convenience, the Hilbert space is divided into two subspaces: stable hydrogen bond subspace and unstable hydrogen bond subspace. Stable subspace includes states $|-1\rangle|0\rangle|1\rangle$ and $|-1\rangle|0\rangle|0\rangle$, and unstable subspace includes the rest. The stable subspace can be defined as follows
$|\rm{O}\cdots\rm{H}\rangle=c^{\prime}|-1\rangle|0\rangle|1\rangle+c^{\prime\prime}|-1\rangle|0\rangle|0\rangle$ (24)
@@
The physical determination of the time of hydrogen bond formation involves a combination of experimental techniques such as ultrafast infrared spectroscopy [51], which determines the time between $0.5ps$ and $1ps$ ($10^{-12}s$). Therefore, in all simulations, two reference parameters are proposed: $g=2\times 10^{-3}\rm{eV},\ \gamma=5\times 10^{-3}\rm{eV}$. All interaction strengths $g_{dist}$ and $g_{prot}$ below are expressed in terms of reference parameter $g$ and all dissipation intensities $\gamma_{isol}^{out}$, $\gamma_{bond}^{out}$ and $\gamma_{phn}^{out}$ below are expressed in terms of reference parameter $\gamma$. Especially, inflow intensities $\gamma_{isol}^{in}$, $\gamma_{bond}^{in}$ and $\gamma_{phn}^{in}$ are not used directly in simulations, but ratios $\mu_{isol}$, $\mu_{bond}$ and $\mu_{phn}$ (inflow intensity divided by dissipation intensity), defined in Eq. (7), are used instead. Except for the above mentioned parameters which will change during the simulation, the remaining parameters are all fixed. Especially, $\hbar\omega_{prot}+\hbar\omega_{dist}=E_{H-bond}$, $\Delta t=0.01\tau$.
Various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model are shown as below five subsections. The supercomputer resources [58] is used to obtain the results of simulations.
@@ IV.1 Comparison between unitary and dissipative evolutions
Firstly, in closed system without the dissipation intensities, the unitary evolution with only three basic states ($|0\rangle|0\rangle|1\rangle$, $|0\rangle|1\rangle|0\rangle$ and $|1\rangle|1\rangle|0\rangle$) is obtained. The unitary evolution considering only the interaction between particles and fields is calculated by the Schrödinger equation.
$i\hbar\dot{\rho}=[H,\rho]=H\rho-\rho H$ (25)
For the unitary evolution, the calculation method only needs to adopt the Eq. (21). The initial state is $|0\rangle|0\rangle|1\rangle$ — stretched and ground state with a phonon. The unitary evolution, considering $g_{prot}=g_{dist}=g$, is shown in Fig. 4. The three obtained curves, representing the states $|0\rangle|0\rangle|1\rangle$, $|0\rangle|1\rangle|0\rangle$ and $|1\rangle|1\rangle|0\rangle$, oscillate periodically.

### v3

@@ II.2 States and Hamiltonian
Figure 2: (online color) Quantum states involved in evolution. (a) $\sim$ (g) correspond to seven states for the $\rm{H}_{2}\rm{O}$-related hydrogen bond model: $|2\rangle|1\rangle|0\rangle,\ |1\rangle|1\rangle|0\rangle,\ |0\rangle|1\rangle|0\rangle,\ |0\rangle|0\rangle|1\rangle,\ |0\rangle|0\rangle|0\rangle,\ |-1\rangle|0\rangle|1\rangle,\ |-1\rangle|0\rangle|0\rangle$. Here, brown, red, green, and purple circles represent oxygen atoms, protons with excited states, protons with ground states, and phonons, respectively; solid and dashed black lines stand for ordinary and stretched covalent bonds, respectively; and ion attraction is denoted by gray dotted lines.
@@
$|d\rangle_{dist}|p\rangle_{prot}|n\rangle_{phn},$ (9)
where the first qubit $|d\rangle_{dist}$ is the relative position of two water molecules, which is detailed in Sec. II.1; the second quantum bit $|p\rangle_{prot}$ is the state corresponding to the proton, $p=0$ — proton is in the ground state $|\Phi_{0}^{pr}\rangle$ and $p=1$ — proton is in the excited state $|\Phi_{1}^{pr}\rangle$; the third qubit $|n\rangle_{phn}$ is the number of phonons corresponding to the hydrogen bond. A special case of our model is proposed, where at most one phonon is pumped into the system (at this time, $n_{av}=1$). When a hydrogen bond is formed, a phonon is emitted. When a hydrogen bond is broken, the phonon is absorbed.
@@
Table 1: Quantum states involved in evolution. The evolution involves a total of 7 states. We use the following modifiers, for convenience, to define the different states: “Broken” means the hydrogen bond is broken and two molecules are in a free state; “critical” means the critical point of the hydrogen bond formation and breaking is obtained; “stretched” means the proton of the donor moves toward the receptor due to the ion attraction, and the covalent bond is stretched; “excited” means the proton state is excited; “ground” means the proton state is ground; and “stable” means the hydrogen bond is stable. In particular, a stable state necessitates both stretching and grounding. Besides states $|-1\rangle|0\rangle|1\rangle$ and $|-1\rangle|0\rangle|0\rangle$, the rest of the states can all be defined as “unstable.” In addition, there are two cases for the ground state: with/without a phonon.
$H_{hb}$ is represented by the Hamiltonian of a system of two water molecules connected by a hydrogen bond and has the following form
$H_{hb}=H_{dist}Cond(p=1)+H_{prot}Cond(d=0),$ (10)
where $H_{dist}$ describes the proton tunneling energy and $H_{prot}$ describes the Hamiltonian for transitions. $Cond(p=1)$ and $Cond(d=0)$ are conditional operators. The rotating wave approximation (RWA) is particularly useful in systems where the interactions are resonant or nearly resonant [55]. RWA is taken into account, and the Hamiltonian is described in the following form
@@
    $H_{dist}=\hbar\omega_{dist}\sigma^{{\dagger}}_{dist}\sigma_{dist}+g_{dist}\left(\sigma^{{\dagger}}_{dist}+\sigma_{dist}\right),$ (11)
    where $\omega_{dist}$ is the frequency for tunneling of phonons; $g_{dist}$, which is located on the non-diagonal line of the Hamiltonian, is the strength for tunneling of protons. And definitions of operators $\sigma_{dist}$ and $\sigma^{{\dagger}}_{dist}$ are as follows
    $\sigma_{dist}|1\rangle_{dist}=|0\rangle_{dist},\ \sigma_{dist}^{{\dagger}}|0\rangle_{dist}=|1\rangle_{dist}.$ (12)
@@
    $H_{prot}=\hbar\omega_{phn}a^{{\dagger}}_{phn}a_{phn}+\hbar\omega_{prot}\sigma^{{\dagger}}_{prot}\sigma_{prot}+g_{prot}\left(a_{phn}\sigma^{{\dagger}}_{prot}+a^{{\dagger}}_{phn}\sigma_{prot}\right),$ (13)
    where $\omega_{phn}$ is the phononic mode, $\omega_{prot}$ is the mode for transitions of protons, and $\omega_{phn}=\omega_{prot}$; $g_{prot}$ is the strength for transitions of protons. And definitions of operators $a^{{\dagger}}_{phn}\sigma_{prot}$ and $a_{phn}\sigma_{prot}^{{\dagger}}$ are as follows
    $a^{{\dagger}}_{phn}\sigma_{prot}|1\rangle_{prot}|0\rangle_{phn}=|0\rangle_{prot}|1\rangle_{phn},\ a_{phn}\sigma_{prot}^{{\dagger}}|0\rangle_{prot}|1\rangle_{phn}=|1\rangle_{prot}|0\rangle_{phn}.$ (14)
The complete quantum system ultimately constructed by the total Hamiltonian $H_{hb}$ can be intuitively represented in the form of a network, which is shown in Fig. 3. The interactions, dissipation, and inflows between the states can be intuitively seen in this figure, and the energy wandering of the whole system is determined clearly.
Figure 3: (online color) Network of quantum states and energy wandering. The quantum system consists of several states, between which there are some potential interactions, dissipation, and inflows. In this figure, the dots represent the states, and the edges represent interactions, dissipation, and inflows. Red dots stand for the states corresponding to the formation of a stable hydrogen bond; the green dot represents the state corresponding to the case that the hydrogen bond is broken and two molecules are free; blue dots represent the states corresponding to an unstable hydrogen bond. Black edges stand for the interactions ($g_{dist}$ and $g_{prot}$); purple edges represent the dissipation and inflows. In particular, $\gamma_{isol}$ includes both dissipation $\gamma_{isol}^{out}$ from state $|1\rangle|1\rangle|0\rangle$ to state $|2\rangle|1\rangle|0\rangle$ and inflow $\gamma_{isol}^{in}$ from state $|2\rangle|1\rangle|0\rangle$ to state $|1\rangle|1\rangle|0\rangle$.
@@ II.3 Quantum master equation
$A_{phn}|1\rangle_{phn}=|0\rangle_{phn},\ A_{phn}^{{\dagger}}|0\rangle_{phn}=|1\rangle_{phn}.$ (15)
And $\gamma_{phn}$ is the dissipation intensity for the escape of the phonons from the cavity to the external environment. Similarly, the two quasi-decoherence factors, which are related to the qubit $|d\rangle_{dist}$ of the basic state, are defined as follows
$A_{bond}|0\rangle_{dist}=|-1\rangle_{dist},\ A_{bond}^{{\dagger}}|-1\rangle_{dist}=|0\rangle_{dist}.$ (16)
$A_{isol}|1\rangle_{dist}=|2\rangle_{dist},\ A_{isol}^{{\dagger}}|2\rangle_{dist}=|1\rangle_{dist}.$ (17)
@@
The QME, which is also called Lindbladian, is used to describe the time-dependent evolution of the density matrix of an open quantum system, which allows us to obtain the dependence of the probability of the hydrogen bond formation on temperature. It can be seen as a natural result of the coupling between the system and the environment under the lowest-order perturbation and Markov approximation
$i\hbar\dot{\rho}=[H,\rho]+i\mathcal{L}(\rho),$ (18)
where $\mathcal{L}=L^{in}+L^{out}$. $L^{in}$ describes the inflow process, and $L^{out}$ describes the dissipation process. Then
$L^{in}(\rho)=\gamma^{in}\left(A^{\dagger}\rho A-\frac{1}{2}\left(AA^{\dagger}\rho+\rho AA^{\dagger}\right)\right),\ L^{out}(\rho)=\gamma^{out}\left(A\rho A^{\dagger}-\frac{1}{2}\left(A^{\dagger}A\rho+\rho A^{\dagger}A\right)\right),$ (19)
where $A$ and $A^{\dagger}$ can be replaced by $A_{bond},\ A_{isol},\ A_{phn},\ A_{bond}^{\dagger},\ A_{isol}^{\dagger}$ and $A_{phn}^{\dagger}$. The full form of the operator $\mathcal{L}$ with consideration of all types of inflows and dissipation is shown in Appx. A.
@@ III Numerical method
We use various numerical methods, such as the Euler method, to solve the QME. The iteration consists of three steps. The first step is to calculate the unitary evolution of the density matrix
$\widetilde{\rho}(t+\Delta t)=e^{\frac{-iH\Delta t}{\hbar}}\rho(t)e^{\frac{iH\Delta t}{\hbar}},$ (20)
where $\Delta t$ is the iteration time step, and the second step is to apply the Lindblad superoperators (19) to the density matrix $\rho(t+\Delta t)$ at this moment
@@
The third step is to scale the density matrix to maintain its positive definiteness, hermiticity, normalization, and other properties.
In addition to the Euler method, there is also the more accurate Runge-Kutta method. For quantum systems with a particularly large number of states, the Monte Carlo wave function method [56] or the pure state vector method [57] can achieve higher computational efficiency. The $\rm{H}_{2}\rm{O}$-related hydrogen bond model contains a total of 7 states, so using the Euler method makes it easier for us to study the effects of various interactions and dissipation on evolution.
When solving the QME using the Euler method, the order of magnitude of the time step $\Delta t$ usually depends on the unit of the system’s Hamiltonian and the units of other related parameters. The measurement time, which is around 1 $\mu s$, determines the time step in the optical cavity QED experiments with high-finesse optical resonators [58] or microwave resonators [59]. However, for a single hydrogen bond, the time step reference measurement time of 1 $\mu s$ is not appropriate. Therefore, the characteristic time scale $\tau$ of the simulated dissipative quantum system is adopted, which can be estimated from the system energy and the reduced Planck constant $\tau=\frac{\hbar}{E}$, where $E$ is the energy of the quantum system and the intensity factor $\gamma$ is usually smaller than $E$. When implementing the Euler method numerically, the time step $\Delta t$ should be significantly smaller than the characteristic time scale.
The energy of a single hydrogen bond $E_{H-bond}$ can be estimated from the average energy of hydrogen bonds in water and Avogadro’s constant, which is approximately 0.217655eV0.217655\ eV. Using the joule as the energy unit may cause computational errors due to extremely small parameter values, so $eV$ is used as the energy unit. And the characteristic time scale $\tau$ is about $3.175316\times 10^{-15}\ s$. The detailed calculation process is shown in Appx. B.
@@ IV Simulations and results
For convenience, the Hilbert space is divided into two subspaces: the stable hydrogen bond subspace and the unstable hydrogen bond subspace. The stable subspace includes the states $|-1\rangle|0\rangle|1\rangle$ and $|-1\rangle|0\rangle|0\rangle$, and the unstable subspace includes the rest. The stable subspace can be defined as follows
$|\rm{O}\cdots\rm{H}\rangle=c^{\prime}|-1\rangle|0\rangle|1\rangle+c^{\prime\prime}|-1\rangle|0\rangle|0\rangle,$ (22)
@@
The physical determination of the time of the hydrogen bond formation involves a combination of experimental techniques such as ultrafast infrared spectroscopy [53], which determines the time between $0.5ps$ and $1ps$ ($10^{-12}s$). Therefore, in all simulations, two reference parameters are proposed: $g=2\times 10^{-3}\rm{eV},\ \gamma=5\times 10^{-3}\rm{eV}$. All interaction strengths $g_{dist}$ and $g_{prot}$ below are expressed in terms of reference parameter $g$, and all dissipation intensities $\gamma_{isol}^{out}$, $\gamma_{bond}^{out}$, and $\gamma_{phn}^{out}$ below are expressed in terms of reference parameter $\gamma$. Especially, inflow intensities $\gamma_{isol}^{in}$, $\gamma_{bond}^{in}$, and $\gamma_{phn}^{in}$ are not used directly in simulations, but ratios $\mu_{isol}$, $\mu_{bond}$, and $\mu_{phn}$ (inflow intensity divided by dissipation intensity), defined in Eq. (7), are used instead. Except for the above-mentioned parameters, which will change during the simulation, the remaining parameters are all fixed. Especially, $\hbar\omega_{prot}+\hbar\omega_{dist}=E_{H-bond}$, $\Delta t=0.01\tau$.
Various dynamic aspects of the $\rm{H}_{2}\rm{O}$-related hydrogen bond model are shown in the five subsections below. The supercomputer resources [60] are used to obtain the results of simulations.
@@ IV.1 Comparison between unitary and dissipative evolutions
Firstly, in a closed system without the dissipation intensities, the unitary evolution with only three basic states ($|0\rangle|0\rangle|1\rangle$, $|0\rangle|1\rangle|0\rangle$, and $|1\rangle|1\rangle|0\rangle$) is obtained. The unitary evolution considering only the interaction between particles and fields is calculated by the Schrödinger equation
$i\hbar\dot{\rho}=[H,\rho]=H\rho-\rho H.$ (23)
For the unitary evolution, the calculation method only needs to adopt Eq. (20). The initial state is $|0\rangle|0\rangle|1\rangle$ — a stretched and ground state with a phonon. The unitary evolution, considering $g_{prot}=g_{dist}=g$, is shown in Fig. 4. The three obtained curves, representing the states $|0\rangle|0\rangle|1\rangle$, $|0\rangle|1\rangle|0\rangle$ and $|1\rangle|1\rangle|0\rangle$, oscillate periodically.
@@
The dissipation intensities are taken into consideration to calculate the dissipative evolution of an open system with phonons entering and exiting. The QME is numerically calculated using Eq. (21). In addition to the initial state $|0\rangle|0\rangle|1\rangle$, the initial state $|1\rangle|1\rangle|0\rangle$ — critical state, is also considered. The critical state means the critical point of the hydrogen bond breaking is reached. A comparison between these two initial states is carried out in Fig. 5. As shown in both panels, the periodic oscillations disappear, replaced by the dissipation. The time of the hydrogen bond formation is about $1\rm{ps}$ ($10^{-12}s$) before the probability reaches a stable level. In panel (a), $\rho_{|-1\rangle|0\rangle|0\rangle\langle-1|\langle 0|\langle 0|}>\rho_{|2\rangle|1\rangle|0\rangle\langle 2|\langle 1|\langle 0|}$, thus, the system tends to form a stable hydrogen bond when the initial state is $|0\rangle|0\rangle|1\rangle$. On the contrary, in panel (b), $\rho_{|-1\rangle|0\rangle|0\rangle\langle-1|\langle 0|\langle 0|}<\rho_{|2\rangle|1\rangle|0\rangle\langle 2|\langle 1|\langle 0|}$, the system tends to break the hydrogen bond when the initial state is $|1\rangle|1\rangle|0\rangle$. This is because the state $|0\rangle|0\rangle|1\rangle$ is stretched and ground, and the proton of the donor at this time is at a lower energy level. Thus, it is more inclined to form a stable hydrogen bond. The state $|1\rangle|1\rangle|0\rangle$ means that the critical point of the hydrogen bond formation and breaking is obtained, and the hydrogen bond breaks more easily to release two free water molecules.
IV.2 The effect of interactions on evolution